OpenStudy (anonymous):

Evaluate 6 + (16-11)2 + √81 ÷ 3

OpenStudy (anonymous):

http://openstudy.com/users/subway_surfers

OpenStudy (anonymous):

please show all the steps

OpenStudy (opcode):

Questions: Is it: \[6 + (16-11)^2 + \sqrt{81} ÷ 3\] or \[6 + (16-11)^2 + \sqrt{81 ÷ 3}\]

OpenStudy (anonymous):

subway_surfers

OpenStudy (anonymous):

@subway_surfers

OpenStudy (anonymous):

do u need me or do u have it explained

OpenStudy (anonymous):

dude coul u do what u did in the last qoustuion?

OpenStudy (the_fizicx99):

The square root of 81 is 9 6 + (16-11)2 + 9 ÷ 3 6 + 5^2 + 9 ÷ 3 6 + 25 + 3 = 34

OpenStudy (anonymous):